# A perspective on inelastic light scattering spectroscopy for probing transport of collective acoustic excitations

### v1

@@ Abstract
Understanding and manipulating nanoscale energy transport and conversion processes are essential for diverse applications, ranging from thermoelectrics and energy harvesting to thermal management of microelectronics. While it has long been recognized that acoustic and thermal properties in condensed matters are primarily due to microscopic transport of phonons as quasiparticles, probing thermal acoustic excitations particularly at sub-THz remains a challenge primarily due to limitations in experimental techniques with spatiotemporal resolutions pertinent to probing them. Brillouin light scattering (BLS) and its variant, impulsive stimulated Brillouin scattering (ISS), provide access to these thermal acoustic excitations, enabling measurement of quantities such as acoustic dispersions along with relaxation dynamics occurring in ultrasonic as well as hypersonic frequencies. In this perspective, we provide a brief overview of the operational principles of BLS and ISS, and highlight their applications in probing acoustic, thermal, and magnetic excitations in emerging and low-dimensional materials. We conclude by discussing current challenges and future opportunities for advanced material characterization using Brillouin light scattering spectroscopy techniques.
@@ Introduction
Nanoscale energy transport and conversion processes are essential for a wide range of applications in thermoelectrics, energy harvesting, and heat dissipation in modern devices Chen (2005); Pop (2010); Gaskins et al. (2017); Scott et al. (2018); Shakouri (2011); Mukherjee et al. (2022); Zhang et al. (2022). These processes are primarily facilitated by quasiparticles, among which phonons are dominant heat carriers in crystals Cahill et al. (2003, 2014). On a microscopic scale, lattice heat conduction can be characterized by dispersion relations determined by crystal structure and relaxation dynamics that govern the transport properties of these heat-carrying quasiparticles. It has been demonstrated that long-wavelength, sub-THz thermal acoustic phonons contribute substantially to heat transport in bulk and nanostructured materials Chen (2005), and that access to these modes is essential for understanding and tailoring their thermal and acoustic properties. Diverse scattering spectroscopy techniques have been developed and applied to probe the transport of thermal phonons in condensed matters, among which inelastic neutron scattering (INS) and inelastic X-ray scattering (IXS) are well-established methods for measuring vibrational dispersion relations and lifetimes Delaire et al. (2011); Ferreira et al. (2018); Masciovecchio et al. (1997); Baldi et al. (2010); Sette et al. (1998). However, since INS requires an ample sample volume (>mm3 Piccoli et al. (2007)), it has limited sample coverage, especially for specimens that need to be synthesized as thin films, and kinematic constraints complicate measurements even when sufficient sample volume is available. Moreover, the resolution limitations of IXS restrict access to sub-THz vibrations. Wischnewski et al. (1998); Baldi et al. (2011).
@@
To probe thermal acoustic phonons beyond the spectral resolution of INS and IXS, Brillouin light scattering spectroscopy (BLS) has emerged as an established technique, complementing Raman spectroscopy which allows to probe zone-center optical polarizations Weber and Merlin (2013); Kargar and Balandin (2021). In principle, Brillouin scattering, also known as Brillouin–Mandelstam scattering, involves the scattering of an optical wave by an acoustic wave. The earliest theoretical works on Brillouin scattering were independently presented by Brillouin Brillouin (1922) and Mandelstam Mandelstam (1926) in the early 20th century. The first experimental observation was reported shortly thereafter by Gross Gross (1930a, b). The subsequent introduction of Fabry-Perot (FP) interferometers by Sandercock Sanderocock (1978) became a key component in modern BLS configurations, providing high contrast while maintaining moderate finesse to ensure sufficient transmittance of Brillouin signal for achieving high signal-to-noise ratio. Later, stimulated Brillouin light scattering was developed as a means to efficiently amplify the scattered signal Chiao et al. (1964) whose intensity by the spontaneous process is known to be substantially weak Damzen et al. (2003); Boyd et al. (2008). The scattering process can be stimulated through the strong coupling between the incident light and the material, producing an exponentially amplified signal Eggleton et al. (2019). Building on this, by introducing a pump laser with a pulse duration less than period of oscillation (inverse of acoustic phonon frequencies), the pump pulses impulsively create spatially periodic and time-dependent excitations in the material Yan et al. (1985); Yan and Nelson (1987). The following modulations in the optical properties of the material can be measured using a probe beam. This method, termed impulsive stimulated Brillouin scattering, was demonstrated by Boersch and Eichler in 1967 Boersch and Eichler (1967). We will refer to this process as impulsive stimulated Brillouin scattering (ISS) when the scattering occurs in a stimulated manner by an impulsive light source, whereas we will specify the term Brillouin light scattering when addressing spontaneous or general Brillouin scattering.
@@
Due to its non-contact and non-destructive nature, BLS has been extensively employed for characterizing photoelastic properties of crystalline and glassy materials across a range of phases Létoublon et al. (2016); Mao et al. (2018); Sandercock (1975); Ko et al. (2011); Benedek and Greytak (1965). In addition to elastic measurements, BLS has emerged as a routine tool for probing properties of quasiparticles such as phonons and magnons. In the case of phonons, BLS enables access to low-frequency phonon dispersion, especially in nanostructured solids. For instance, BLS has been used to observe confined phonon modes in GaAs nanowires Kargar et al. (2016), and to resolve gigahertz phonon dispersion relations in patterned holey silicon membranes, thereby providing direct experimental evidence of omnidirectional vibrational band gaps in phononic crystals Florez et al. (2022). Beyond acoustic excitations, BLS can also probe magnons, which are of interest for spintronics applications Demidov et al. (2009); Vogt et al. (2014); Nembach et al. (2015); Holanda et al. (2018); Demidov et al. (2016). BLS has been employed to spatially map magnon propagation in magnonic waveguides Demidov et al. (2007); Vogt et al. (2012); Demidov et al. (2016), to quantify interfacial Dzyaloshinskii-Moriya interaction strength in metal multilayers Nembach et al. (2015); Cho et al. (2015) and, notably, to demonstrate magnon Bose-Einstein condensate at room temperature Demokritov et al. (2006); Borisenko et al. (2020).
@@
Stimulated Brillouin scattering and impulsive stimulated Brillouin scattering extend these capabilities by enhancing the intensity of Brillouin signal. These techniques have been widely applied to study thermal and elastic properties of thin films, suspended membranes, and bulk materials Johnson et al. (2012, 2013); Cuffe et al. (2015); Li et al. (2021), as well as highly anisotropic materials and soft matters Rogers et al. (2000); Krug et al. (2019); Hecksher et al. (2017); Khanolkar et al. (2015). A key feature of impulsive stimulated Brillouin scattering is its tunable accessible length, which is pertinent to probing transport processes across different spatio-temporal scales. This extends measurements of heat transport beyond the diffusive regime, allowing experimental reconstruction of vibrational mean free paths (MFPs) spectra in lattices with various phases by resolving ballistic heat transport Johnson et al. (2013); Cuffe et al. (2015); Robbins et al. (2019); Kim et al. (2021). Additionally, beyond mean free path spectroscopy, it has been applied to detect second sound in graphite at markedly high temperatures by accessing phonon hydrodynamic transport regime Huberman et al. (2019); Ding et al. (2022). Another variant of ISS is spin grating spectroscopy, which has been applied to probe electronic spin dynamics and valley polarization dynamics: discovery of spin-Coulomb drag in GaAs quantum wells; observation of the persistent spin helix in semiconductor quantum wells; measurement of valley depolarization dynamics in low-dimensional materials Weber et al. (2005); Koralek et al. (2009); Mahmood et al. (2018).
@@ I  Working principles and instrumentation of BLS
Unlike other inelastic scattering spectroscopies such as inelastic X-ray scattering, which are limited to resolve excitations with frequencies above a few THz, Brillouin scattering offers access to acoustic excitations with frequencies ranging from sub-gigahertz up to several hundred gigahertz, as shown in Fig. 1(a). In this section, we provide a brief introduction on the working principle of Brillouin scattering spectroscopy techniques, and outline modern experimental configurations that enable such spectral resolution in these photonic approaches.
@@ I.1  Stokes and anti-Stokes scattering
Typical BLS spectra exhibit several intensity peaks as shown in Fig. 1(b): a peak with its central frequency identical to that of the incident optical wave as a result of elastic Rayleigh scattering, along with other frequency-shifted (inelastic) peaks demonstrating Stokes shifts. The downshifted component is due to Brillouin scattering in which light is scattered by an interaction with acoustic waves, namely phonons, in the medium. In this process, called the Stokes process, phonons are created during the light-matter interaction, resulting in a lower frequency of the scattered wave. In contrast, the upshifted peak at higher frequency corresponds to the anti-Stokes scattering where phonons are annihilated during the interaction, causing a frequency increase in the scattered light.
The inelastic scattering events should obey energy and momentum conservation between the incident and scattered lights with nearly identical wavevectors due to small momentum transfer compared to the optical wavevectors $\mathbf{k}$. The frequency shift of measured lines, which is identical to those of phonons with momentum $\mathbf{q}$ is expressed as Boyd et al. (2008)
$\nu=\pm\frac{2nv_{s}}{\lambda_{i}}\sin\left({\theta}/{2}\right)$ (1)
@@
where $\lambda_{i}$ is wavelength of incident optical wave, $n$ is the refractive index, $v_{s}$ is acoustic velocity, and $\theta$ is scattering angle. As can be seen in Eq. 1, it is apparent that the magnitude of the frequency shift is maximum (minimum) at scattering angle of 180°(0°). This indicates that backscattering geometry ($\theta=180\degree$) provides access to the maximum detectable phonon wavevector along with the corresponding frequency, while spectrum of forward scattering will be dominated by Rayleigh scattering due to the smaller momentum transfer. The frequency shift in Brillouin scattering can also be interpreted as a Doppler effect between light and the acoustic wave. More precisely, the redshift observed in Stokes scattering is analogous to light scattering from a receding acoustic wave in the material, while the blueshift of the scattered wave in anti-Stokes scattering corresponds to scattering from an approaching acoustic wave Boyd et al. (2008); Benedek and Fritsch (1966).
@@
An additional physical property can be extracted from analyzing the peak widths of BLS spectra. As shown in the peaks of Fig. 1(b), shifted peaks manifest a finite width due to acoustic attenuation. The linewidths of the peaks are directly related to the phonon lifetimes, where phonon lifetime $\tau_{p}$ can be determined using full width at half maximum (FWHM) of the Brillouin peak using $\Delta\nu=1/\Delta(2\pi\,\tau_{p})=\Gamma{\lvert\mathbf{q}\rvert}^{2}/2\pi$. Here, $\Delta\nu$ is the FWHM and $\Gamma$ is the attenuation coefficient Berne and Pecora (2000). Noting that the momentum transfer is $\lvert\mathbf{q}\rvert=2\lvert\mathbf{k}\rvert\sin(\theta/2)$, lifetime is determined as follows Boyd et al. (2008).
@@
From Eq. 2, the width of Brillouin peaks directly reflects the lifetimes of phonons which participate in the scattering process, and provides insight into the transport properties of low-energy phonons in the material along with the physical properties such as elastic constants. For example, in liquids, sub-ns vibrational lifetimes at a few gigahertz were measured in toluene and methylene chloride, and the results reflect the viscosity coefficients of such media Shapiro et al. (1966). The vibrational lifetimes in several glasses were estimated to be in the range of 2-4 ns at hypersonic frequencies, falling between the values of organic liquids and crystalline quartz, where the latter exhibits a phonon lifetime of $\sim 5\,\mathrm{ns}$ Pohl and Kaiser (1970). Given that lifetimes of acoustic vibrations in glasses lie between those of crystals and liquids, it indicates that vibrational lifetimes depend on the underlying lattice structure and phase of a material.

### v2

@@ Abstract
Understanding and manipulating nanoscale energy transport and conversion processes are essential for diverse applications, ranging from thermoelectrics and energy harvesting to thermal management of microelectronics. While it has long been recognized that acoustic and thermal properties in condensed matters are primarily due to microscopic transport of phonons as quasiparticles, probing thermal acoustic excitations particularly at sub-THz remains a challenge primarily due to limitations in experimental techniques with spatiotemporal resolutions pertinent to probing them. Brillouin light scattering (BLS) and its variant, impulsive stimulated scattering (ISS), provide access to these thermal acoustic excitations, enabling measurement of quantities such as acoustic dispersions along with relaxation dynamics occurring in ultrasonic as well as hypersonic frequencies. In this perspective, we provide a brief overview of the operational principles of BLS and ISS, and highlight their applications in probing acoustic, thermal, and magnetic excitations in emerging and low-dimensional materials. We conclude by discussing current challenges and future opportunities for advanced material characterization using Brillouin light scattering spectroscopy techniques.
@@ Introduction
Nanoscale energy transport and conversion processes are essential for a wide range of applications in thermoelectrics, energy harvesting, and heat dissipation in modern devices Gang_2005_txtbk; Pop_2010_NanoR; Gaskins_2017_ElecSoc_highkdi_review; Scott_2018_APLmat_highkdi; Shakouri_2011_AnnuRev_thermoelectric1; Mukherjee_2022_JMCC_thermoelectric2; Zhang_2022_NatElec_thermoelectric3. These processes are primarily facilitated by quasiparticles, among which phonons are dominant heat carriers in crystals Cahill_2003_JAP_nanoheat; Cahill_2014_APR_nanoheat2. On a microscopic scale, lattice heat conduction can be characterized by dispersion relations determined by crystal structure and relaxation dynamics that govern the transport properties of these heat-carrying quasiparticles. It has been demonstrated that long-wavelength, sub-THz thermal acoustic phonons contribute substantially to heat transport in bulk and nanostructured materials Gang_2005_txtbk, and that access to these modes is essential for understanding and tailoring their thermal and acoustic properties. Diverse scattering spectroscopy techniques have been developed and applied to probe the transport of thermal phonons in condensed matters, among which inelastic neutron scattering (INS) and inelastic X-ray scattering (IXS) are well-established methods for measuring vibrational dispersion relations and lifetimes Delaire_2011_Natmat_PbTe; Ferreira_2018_PRL; Masciovecchio_prb_1997; baldi_prl_2010; sette_science_1998. However, both techniques face constraints arising from instrumental limitations and scattering kinematics. In IXS, its limited energy resolution restricts access to sub-THz vibrations Wischnewski_1998_PRB; Baldi_2011_JNonCS, and achieving high resolution requires high brilliance synchrotron sources, restricting its use to large-scale facilities Baron_2016_intro. INS, on the other hand, can provide enhanced energy resolution and is less invasive than IXS owing to high penetration depth of neutrons, which deposit relatively small energy in the sample, whereas X-rays typically have shallow penetration depth and can induce radiation damage Butler_2022_book. Nevertheless, since INS requires an ample sample volume (>mm3 Piccoli_2007_INS), it has limited sample coverage, especially for specimens that need to be synthesized as thin films, and kinematic constraints complicate measurements even when sufficient sample volume is available.
@@
To probe thermal acoustic phonons beyond the spectral resolution of INS and IXS, Brillouin light scattering spectroscopy (BLS) has emerged as an established technique, complementing Raman spectroscopy which allows to probe zone-center optical polarizations weber_2013_raman; Kargar_2021_natphot. In principle, Brillouin scattering, also known as Brillouin–Mandelstam scattering, involves the scattering of an optical wave by an acoustic wave. The earliest theoretical works on Brillouin scattering were independently presented by Brillouin brillouin1922 and Mandelstam mandelstam1926 in the early 20th century. The first experimental observation was reported shortly thereafter by Gross gross1930; gross1930_2. The subsequent introduction of Fabry-Perot (FP) interferometers by Sandercock sandercock_1978_SSC became a key component in modern BLS configurations, providing high contrast while maintaining moderate finesse to ensure sufficient transmittance of Brillouin signal for achieving high signal-to-noise ratio. BLS has since become widely accessible owing to their compact optical setup and automated FP interferometers, enabling routine measurements of acoustic excitations with sub-GHz spectral resolution in standard laboratory settings hillebrands_revsciinst_1999; Bencivenga_revsciinst_2012. Later, stimulated light scattering was developed as a means to efficiently amplify the scattered signal Chiao_1964_PRL whose intensity by the spontaneous process is known to be substantially weak damzen_2003_stimulated; Boyd_2008_textbook. The scattering process can be stimulated through the strong coupling between the incident light and the material, producing an exponentially amplified signal Eggleton_NatPhoto_2019. Building on this, by introducing a pump laser with a pulse duration less than period of oscillation (inverse of acoustic phonon frequencies), the pump pulses impulsively create spatially periodic and time-dependent excitations in the material Yan_1985_JCP_ISS1; Yan_1987_JCP_ISS2. The following modulations in the optical properties of the material can be measured using a probe beam. This method, termed impulsive stimulated scattering, was first demonstrated by Boersch and Eichler in 1967 Boersch_1967_firstISBS. We will refer to this process as impulsive stimulated scattering (ISS) when the scattering occurs in a stimulated manner by an impulsive light source, whereas we will specify the term Brillouin light scattering (BLS) when addressing spontaneous or general Brillouin scattering.
@@
Due to its non-contact and non-destructive nature, BLS has been extensively employed for characterizing photoelastic properties of crystalline and glassy materials across a range of phases letoublon2016elastic_BLSapplication1; mao2018solids; sandercock2007some; ko2011precursor; benedek_ieee_1965_liquid. In theory, spontaneous Brillouin scattering involves no optical absorption, therefore impacts of heating in BLS are typically insufficient to induce sample damage Nikolic_2019_biomedoptexp. Moreover, unlike conventional static methods based on finite-deformation stress-strain measurements, which typically yield a few elastic constants under stringent orientation conditions Koski_2013_natmat; Hearmon_1946_revmodphys, BLS enables evaluation of the full elastic tensor in anisotropic materials and oriented fibers Koski_2013_natmat; Elsayad_2020_cellulose. In addition to elastic measurements, BLS has emerged as a routine tool for probing properties of quasiparticles such as phonons and magnons. In the case of phonons, BLS enables access to low-frequency phonon dispersion, especially in nanostructured solids. For instance, BLS has been used to observe confined phonon modes in GaAs nanowires Kargar_2016_Natcom_BLSGaAs, and to resolve gigahertz phonon dispersion relations in patterned holey silicon membranes, thereby providing direct experimental evidence of omnidirectional vibrational band gaps in phononic crystals Florez_2022_NatNanotech. Beyond acoustic excitations, BLS can also probe magnons, which are of interest for spintronics applications Demidov_2009_prl; Vogt_2014_natcomm; Nembach_natphys_2015; Holanda_2018_natphys; Demidov_2016_natcomm. BLS has been employed to spatially map magnon propagation in magnonic waveguides demidov_apl_2007; vogt_apl_2012; Demidov_2016_natcomm, to quantify interfacial Dzyaloshinskii-Moriya interaction strength in metal multilayers Nembach_natphys_2015; cho_natcommun_2015 and, notably, to demonstrate magnon Bose-Einstein condensate at room temperature Demokritov_2006_Nat_BoseEinstein; Borisenko_2020_natcomm. However, despite these capabilities, BLS suffers from an inherently weak scattering signal, resulting in long acquisition times and low signal-to-noise ratio (SNR) Kabakova_natrev_2024; li_photonres_2024; Bilenca_2024_jphysphot.
ISS can access dynamical responses with temporal resolution below nanosecond and has tunable accessible length scales, providing enhanced signal intensity with an additional laser beam, though optical wave mixing geometry is required Choudhry_2021_JAP. Here, ISS can be categorized into impulsive stimulated Brillouin scattering (ISBS), which arises from electrostriction in transparent materials, and impulsive stimulated thermal scattering (ISTS), which occurs in optically absorbing media. While ISBS has been traditionally used to measure viscoelastic properties of liquids yan_jchemphys_1988; Silence_1992_jchemphys; Le_2024_jphysphot, recent development of ISBS microscopy has shifted its focus toward elastic characterizations of biological specimens Ballmann_2017_optica; krug2019impulsive; li_photonres_2024. Owing to enhanced intensity of Brillouin signal generated through impulsive stimulated processes, ISBS offers improved spectral resolution, shorter acquisition times, and higher SNR compared to BLS Prevedel_2019_NatMethods_BLSmicroscopy. However, it is more invasive due to the increased optical power exposure, which can lead to photodamage in sensitive and fragile specimens Chow_2023_optexp. This has motivated ongoing efforts to mitigate laser-induced damage while enhancing SNR li_photonres_2024. Another active area of research in ISBS focuses on pushing the limit of spatial resolution, as its current resolution remains limited to several tens to hundreds of micrometers Zhang_2021_natprot.
@@
ISTS has been widely applied to study thermal and elastic properties of thin films, suspended membranes, and bulk materials Johnson_2012_JAP; Johnson_2013_PRL_TGsignals; cuffe2015reconstructing; li2021remarkably, as well as highly anisotropic materials and soft matters rogers2000optical; hecksher2017toward; khanolkar_apl_2015. A key feature of impulsive stimulated Brillouin scattering is its tunable accessible length, which is pertinent to probing transport processes across different spatio-temporal scales. This extends measurements of heat transport beyond the diffusive regime, allowing experimental reconstruction of vibrational mean free paths (MFPs) spectra in lattices with various phases by resolving ballistic heat transport Johnson_2013_PRL_TGsignals; cuffe2015reconstructing; Andrew_2019_PNAS_DrawRatio; Kim_2021_PRM_aSi_TG. Additionally, beyond mean free path spectroscopy, it has been applied to detect second sound in graphite at markedly high temperatures by accessing phonon hydrodynamic transport regime Huberman_2019_Science; ZDing_2022_NatComm. Another variant of ISS is spin grating spectroscopy, which has been applied to probe electronic spin dynamics and valley polarization dynamics: discovery of spin-Coulomb drag in GaAs quantum wells; observation of the persistent spin helix in semiconductor quantum wells; measurement of valley depolarization dynamics in low-dimensional materials weber_nature_2005; koralek2009emergence; Mahmood_2018_NanoLetters_ISBSapp2.
The interpretation of BLS and ISS measurements require subsequent signal processing and data analysis. For post-processing in BLS, the measured BLS spectrum must be corrected for instrumental broadening arising from the laser linewidth and the finite finesse of FP interferometers Hwa_joptsoc_1989. The observed Brillouin linewidth therefore represents a convolution of the intrinsic phonon linewidth with the instrument response function Bouvet_natphot_2025. The intrinsic linewidth, hence the phonon lifetime, is obtained by deconvolving this response through analytical or numerical fitting approaches  Hwa_joptsoc_1989; Vanderwal_1981_optcommun. In ISBS, the impulsive excitation of the medium by crossed pump pulses launches damped standing-wave oscillations Yan_1987_JCP_ISS2. Traditionally, the resulting signal is analyzed by nonlinear least-squares fitting to a damped sinusoidal function to extract the acoustic frequency and attenuation rate. Subsequently, viscoelastic properties are determined using relaxation models, such as Debye or Kohlrausch-Williams-Watts (KWW) models Yan_1987_JCP_ISS2; Duggal_1991_jchemphys. While ISBS and ISTS measurements share similar geometry, ISTS involves temperature rise due to optical absorption, whereas ISBS is excited by electrostrictive forces Torchinsky_JCP_2009. In ISTS, signals from the photoexcitation process by the impulsive laser beam primarily comprise electronic and thermal decay responses of the absorbing material. Rather than directly deconvoluting the signal, the fitting function is convolved with the impulse response, which contains the instrument’s response Andrew_2019_PNAS_DrawRatio; Kim_Natcomm_2022. The electronic and thermal relaxation dynamics are then obtained by extracting the decay time constants from the fits of the measured signals, and further analysis of phonon dynamics can be conducted based on grating period dependency.
@@ I  Working principles and instrumentation of BLS
Unlike other inelastic scattering spectroscopies such as inelastic X-ray scattering, which are limited to resolve excitations with frequencies above a few THz, Brillouin scattering offers access to acoustic excitations with frequencies ranging from sub-gigahertz up to several hundreds gigahertz, as shown in Fig. 1(a). In this section, we provide a brief introduction on the working principle of Brillouin scattering spectroscopy techniques, and outline modern experimental configurations that enable such spectral resolution in these photonic approaches.
Figure 1: (a) Map showing the accessible frequency and momentum transfer ranges of different inelastic scattering spectroscopy techniques. Shaded regions indicate typical accessible windows for each technique, and the Brillouin scattering shows the appropriate range to probe excitations with relatively low frequency and momentum. Adapted from Ref. Bencivenga_APX_2023 with permission under $C$ BY license. (b) Representative spectra for Brillouin scattering, showing inelastic Stokes (downshifted) and anti-Stokes (upshifted) peaks displaced from the central elastic Rayleigh scattering peak by Brillouin shift. The shifted peaks exhibit relatively low intensity compared to Rayleigh scattering, and the spectral separation is determined by the wavelength and angle of incident light, together with the refractive index and acoustic velocity of a material. The dotted lines represent the measurement range of BLS, Raman, and low-wavenumber Raman (LWNR) spectroscopy. Reprinted with permission from Karger, F. et al., Nature Photon. 15, 720-731 (2021). Copyright 2021 Springer Nature. (c) Measured mean free paths versus frequency for thermal acoustic vibrations in vitreous silica from IXS (triangles, Ref. baldi_prl_2010; Masciovecchio_prb_1997), INS (open circles, Ref. Wischnewski_prb_1998), BLS (5-pointed stars, Ref. Masciovecchi_prl_2006; Benassi_prb_2005; Vacher_prb_2006; Levelut_prb_2006).
@@ I.1  Stokes and anti-Stokes scattering
Typical BLS spectra exhibit several intensity peaks as shown in Fig. 1(b): a central peak occurring at zero frequency is due to elastic Rayleigh scattering, along with other frequency-shifted inelastic peaks. The downshifted peak is due to Brillouin scattering in which light is scattered by interaction with acoustic waves, namely phonons, in the medium. In this process, called the Stokes process, phonons are created during the light-matter interaction, resulting in a lower frequency of the scattered wave. In contrast, the upshifted peak at higher frequency corresponds to the anti-Stokes scattering where phonons are annihilated during the interaction, causing a frequency increase in the scattered light. For both Stokes and anti-Stokes Brillouin peaks, the frequency shifts relative to elastic peak correspond to the frequency of excitations (e.g., phonons) with wavevectors following the dispersion relation, while the widths of the inelastic peaks reflect the phonon lifetimes. The phonon wavevector can be calculated from the energy and momentum conservation of Brillouin scattering and the phonon lifetime can be determined from Lorentzian fitting of the Brillouin peaks.
During inelastic light scattering processes, the incident and scattered light with wavevectors $\mathbf{k}$ and $\mathbf{k^{\prime}}$, respectively, should satisfy the conservation of momentum $\mathbf{k^{\prime}}=\mathbf{k}\pm\mathbf{q}$, where $\mathbf{q}$ is the momentum transfer, identical to the wavevector of an acoustic wave with frequency $\nu$. Here, the momentum transfer is substantially smaller compared to the incident and scattered light wavevectors, and thus can be considered negligible. Then, the momentum transfer becomes $\left|\mathbf{q}\right|=2\left|\mathbf{k}\right|\sin{\left({\theta}/{2}\right)}$, where $\theta$ is the scattering angle. Finally, following the dispersion relation $\nu={\left|\mathbf{q}\right|v_{s}}/{2\pi}$, the frequency shift $\nu$ of the measured lines, which is identical to the frequency of the acoustic wave with wavevector $\mathbf{q}$, is expressed as Boyd_2008_textbook
$\nu=\frac{2nv_{s}}{\lambda_{i}}\sin\left({\theta}/{2}\right)$ (1)
@@
where $\lambda_{i}$ is the wavelength of the incident optical wave, $n$ is the refractive index, and $v_{s}$ is the acoustic velocity. As can be seen in Eq. 1, it is apparent that the magnitude of the frequency shift is maximum (minimum) at scattering angle of 180°(0°). This indicates that backscattering geometry ($\theta=180\degree$) provides access to the maximum detectable phonon wavevector along with the corresponding frequency, while spectrum of forward scattering will be dominated by Rayleigh scattering due to the smaller momentum transfer. The frequency shift in Brillouin scattering can also be interpreted as a Doppler effect between light and the acoustic wave. More precisely, the redshift observed in Stokes scattering is analogous to light scattering from a receding acoustic wave in the material, while the blueshift of the scattered wave in anti-Stokes scattering corresponds to scattering from an approaching acoustic wave Boyd_2008_textbook; Benedek_1966_PhysRev.
@@
An additional physical property can be extracted from analyzing the peak widths of BLS spectra. As shown in the peaks of Fig. 1(b), shifted peaks manifest a finite width due to acoustic attenuation. The linewidths of the peaks are directly related to the phonon lifetimes, where phonon lifetime $\tau_{p}$ can be determined using full width at half maximum (FWHM) of the Brillouin peak using $\Delta\nu=1/2\pi\,\tau_{p}=\Gamma{\lvert\mathbf{q}\rvert}^{2}/2\pi$. Here, $\Delta\nu$ is the FWHM and $\Gamma$ is the attenuation coefficient berne_2000_text. Noting that the momentum transfer is $\lvert\mathbf{q}\rvert=2\lvert\mathbf{k}\rvert\sin(\theta/2)$, lifetime is determined as follows Boyd_2008_textbook.
@@
From Eq. 2, the width of Brillouin peaks directly reflects the lifetimes of phonons which participate in the scattering process, and provides insight into the transport properties of low-energy phonons in the material along with the physical properties such as elastic constants. For example, in liquids, sub-ns vibrational lifetimes at a few gigahertz were measured in toluene and methylene chloride, and the results reflect the viscosity coefficients of such media shapiro_ieee_1966. The vibrational lifetimes in several glasses were estimated to be in the range of 2-4 ns at hypersonic frequencies, falling between the values of organic liquids and crystalline quartz, where the latter exhibits a phonon lifetime of $\sim 5\,\mathrm{ns}$ Kaiser_PRB_1970. Given that lifetimes of acoustic vibrations in glasses lie between those of crystals and liquids, it indicates that vibrational lifetimes depend on the underlying lattice structure and phase of a material. Furthermore, the mean free paths of thermal acoustic excitations can be obtained from the measured lifetimes and velocity determined by BLS. Figure 1(c) illustrates the measured mean free paths versus frequency for vitreous silica obtained from BLS, INS, and IXS, highlighting how BLS bridges the low-frequency regime inaccessible using conventional inelastic scattering spectroscopy techniques.